OpenStudy (anonymous):

2 2/3 of what number is 6 1/7 and how?

OpenStudy (anonymous):

This can be written as: 2 2/3 x = 6 1/7 To find x, divide both sides by 2 2/3 It might be easier to see if you convert to improper fractions: 2 2/3 = 6/3+2/3 = 8/3 6 1/7 = 42/7+1/7 = 43/7 "what number" = "x" (this is the number you're looking for) So now we have 8/3x = 43/7 To find x, divide both sides by 8/3 Your answer this way will be an improper fraction, so you'll have to reduce it.